# On the Effectiveness of ASR Representations in Real-world Noisy Speech Emotion Recognition

### v3

@@ V-E Impact of ASR Representations in Speech and Text Modalities on NSER
To investigate the impact of ASR representations in speech and text modalities on NSER, we conduct a comparative analysis, contrasting different text inputs for BERT in the text modality with various ASR representations in the speech modality, as detailed in Table IV.
@@
In the text modality, BERT applied to transcripts consistently outperforms BERT on Whisper-processed text, with notable performance disparities, particularly on the MELD dataset. However, across the MELD, IEMOCAP, and IEMOCAP-Environment datasets, the speech modality demonstrates superior performance over the text modality, highlighting the added value of ASR representations for NSER.
In the speech modality, the Encoder+Decoder consistently yields the best results, outperforming the individual contributions of the encoder or decoder. This indicates the synergistic potential of integrating comprehensive ASR strategies for NSER. Specifically, on the MELD dataset, the Encoder+Decoder achieves improvements of 5.77 in UAR and 3.37 in F1 compared with BERT on transcripts. Even the Encoder and Decoder individually provide substantial gains, with the Encoder alone improving UAR and F1 by 3.18 and 0.96, and the Decoder contributing gains of 3.19 and 2.05 in UAR and F1, respectively.
@@
On the IEMOCAP dataset, similar trends are observed. The Encoder+Decoder demonstrates a significant advantage over BERT on transcripts, with improvements of 6.41 in UAR and 6.94 in F1. The Encoder independently contributes gains of 4.93 in UAR and 4.51 in F1, whereas the Decoder provides improvements of 2.76 in UAR and 2.80 in F1.
In the Environment subset, the Encoder+Decoder outperforms BERT on transcripts by 2.91 in UAR and 2.83 in F1. In the Human subset, the most challenging noise condition, the Encoder+Decoder achieves slightly lower performance than BERT on transcripts, with decreases of 1.3 in UAR and 1.2 in F1.
Overall, these findings emphasize the overall superiority of the speech modality compared with the text modality in NSER tasks, especially under environmental and mixed noise conditions.
V-F Robustness of ASR Representations in Scenarios Involving Cross-lingual NSER
@@
To evaluate the robustness of ASR representations in cross-lingual NSER tasks, we conducted experiments in which neither the ASR nor SSL models were exposed to the target language during training. On the other hand, a monolingual ASR model was used as a baseline for comparison. Whisper served as the ASR model, generating both encoder and decoder representations. Table V summarizes the performance of SSL models, Mel-Spectrogram, and Whisper-based ASR representations. The abbreviation “FZ” indicates frozen models, “FT” denotes fine-tuned models, “EN” refers to Whisper trained on the English dataset, and “CN” refers to Whisper trained on the Mandarin Chinese dataset.
@@
SER | 53.53 | 52.91 | ||
| Waveform | Wav2Vec 2.0 (FZ) + SER | 22.00 | 18.79 | ||
| Hubert (FZ) + SER | 26.49 | 22.88 | |||
| WavLM (FZ) + SER | 21.47 | 18.80 | |||
| Wav2Vec 2.0 (FT) + SER | 52.44 | 52.39 | |||
| Hubert (FT) + SER | 51.44 | 50.88 | |||
| WavLM (FT) + SER | 53.89 | 52.17 | |||
| Whisper-Encoder(EN) + SER | 52.90 | 52.84 | |||
| Whisper-Decoder(EN) + SER | 55.30 | 54.26 | |||
| Whisper-Encoder+Decoder(EN) + SER | 55.90 | 54.52 | |||
| Whisper-Encoder(CN) + SER | 56.43 | 55.25 | |||
| Whisper-Decoder(CN) + SER | 57.41 | 56.71 | |||
| Whisper-Encoder+Decoder(CN) + SER | 58.54 | 58.13 |
@@
To further clarify the performance impact of language mismatch, we compared Whisper-EN and Whisper-CN representations under cross-lingual and monolingual scenarios, respectively. Compared with Whisper-CN, Whisper-EN shows in UAR and F1 declines across all components. For the Encoder, there is a decrease of 3.53 in UAR and 2.41 in F1, whereas the Decoder experiences reductions of 2.11 in UAR and 2.45 in F1. The Encoder+Decoder shows the largest decline, with 2.64 lower UAR and 3.61 lower F1, highlighting the compounded challenges of language mismatch in cross-lingual settings.
Overall, Whisper-based ASR representations consistently surpass both SSL and Mel-Spectrogram baselines in cross-lingual and monolingual contexts, demonstrating remarkable robustness against language mismatch and noise.
@@ VI Conclusions and Future Work
In this paper, we presented a novel approach to NSER by leveraging ASR models as robust feature extractors in noisy environments. Our investigation demonstrated the effectiveness of this method in addressing the limitations of conventional NSER techniques when faced with real-world noise challenges. By utilizing intermediate layer representations from ASR models, we captured emotional speech features with greater accuracy, resulting in superior NSER performance compared with conventional noise reduction methods and SSL approaches.
@@
Looking ahead, future work could further refine our approach by incorporating advanced ASR models or exploring alternative feature extraction techniques to boost NSER performance. Expanding evaluations to a broader range of datasets and exploring the method’s robustness across different languages and cultural contexts will enhance our understanding and practical application of SER.

### v4

@@ V-F Impact of ASR Representations in Speech and Text Modalities on NSER
To investigate the impact of ASR representations in speech and text modalities on NSER, we conduct a comparative analysis, contrasting different text inputs for BERT in the text modality with various ASR representations in the speech modality, as detailed in Table VI.
@@
In the text modality, BERT applied to transcripts consistently outperforms BERT on Whisper-generated text, with the largest margin observed on the IEMOCAP-Human subset. However, across most datasets, the speech modality demonstrates superior performance compared with the text modality, except for the IEMOCAP-Human subset, highlighting the added value of ASR representations for NSER.
For the MELD dataset, the Proposed method achieves 48.95% UAR and 48.87% F1, improving over BERT (Transcripts) by 5.77% in UAR and 3.37% in F1. The Encoder alone improves by 3.53% in UAR and 2.51% in F1, while the Decoder adds 3.19% in UAR and 2.05% in F1.
@@
For the IEMOCAP dataset, the Proposed method reaches 76.50% UAR and 76.53% F1, yielding improvements of 6.41% in UAR and 6.94% in F1 over BERT (Transcripts). The Encoder contributes 4.93% in UAR and 4.51% in F1, while the Decoder provides 2.66% in UAR and 2.80% in F1.
For the IEMOCAP-Human subset, BERT (Transcripts) achieves 70.09% UAR and 69.59% F1, which remains higher than the speech-based alternatives. The Proposed method obtains 68.79% UAR and 68.39% F1, falling short by 1.30% in UAR and 1.20% in F1 compared with BERT (Transcripts). The Encoder lags by 3.05% in UAR and 3.23% in F1, while the Decoder is lower by 3.10% in UAR and 2.98% in F1.
For the IEMOCAP-Environment (Park) subset, the Proposed method reaches 73.00% UAR and 72.42% F1, surpassing BERT (Transcripts) by 2.91% in UAR and 2.83% in F1. The Encoder provides smaller gains of 0.47% in UAR and a slight decrease of 0.16% in F1, while the Decoder remains lower by 0.37% in UAR and 1.20% in F1.
For the IEMOCAP-Environment (Traffic) subset, the Proposed method achieves 74.39% UAR and 74.04% F1, outperforming BERT (Transcripts) by 4.3% in UAR and 4.45% in F1. The Encoder independently improves by 2.12% in UAR and 2.28% in F1, while the Decoder contributes 1.81% in UAR but falls behind by 0.38% in F1.
Overall, these findings show that ASR-based speech representations generally outperform text-based representations for NSER, particularly under environmental and mixed noise conditions where robustness is critical. However, in the Human subset, BERT on clean transcripts remains slightly stronger, suggesting that text modality can still provide advantages when high-quality transcriptions are available. This highlights the complementary roles of speech and text modalities in NSER.
V-G Robustness of ASR Representations in Scenarios Involving Cross-lingual NSER
@@
To evaluate the robustness of ASR representations in cross-lingual NSER tasks, we conducted experiments where neither the ASR nor SSL models were exposed to the target language during training. A monolingual ASR model was used as a baseline. Whisper served as the ASR model, providing both encoder and decoder representations. Table VII summarizes the performance of SSL models, Mel-Spectrogram, and Whisper-based ASR representations.
@@
BiGRU | 53.53 | 52.91 | ||
| Waveform | Wav2Vec 2.0 (FZ) | 22.00 | 18.79 | ||
| HuBERT (FZ) | 26.49 | 22.88 | |||
| WavLM (FZ) | 21.47 | 18.80 | |||
| Wav2Vec 2.0 (FT) | 52.44 | 52.39 | |||
| HuBERT (FT) | 51.44 | 50.88 | |||
| WavLM (FT) | 53.89 | 52.17 | |||
| Whisper-Enc(EN) | 52.90 | 52.84 | |||
| Whisper-Dec(EN) | 55.30 | 54.26 | |||
| Whisper-Enc+Dec(EN) | 55.90 | 54.52 | |||
| Whisper-Enc(CN) | 56.43 | 55.25 | |||
| Whisper-Dec(CN) | 57.41 | 56.71 | |||
| Whisper-Enc+Dec(CN) | 58.54 | 58.13 |
@@
To further clarify the impact of language mismatch, we compared Whisper-EN and Whisper-CN representations under cross-lingual and monolingual scenarios. Relative to Whisper-CN, Whisper-EN shows declines in both UAR and F1 across all components. For the Encoder, there is a decrease of 3.53 in UAR and 2.41 in F1, whereas the Decoder experiences reductions of 2.11 in UAR and 2.45 in F1. The Encoder+Decoder shows the largest decline, with 2.64 lower UAR and 3.61 lower F1, highlighting the compounded challenges of language mismatch in cross-lingual settings.
Overall, Whisper-based ASR representations consistently surpass both SSL and Mel-Spectrogram baselines in cross-lingual and monolingual contexts, demonstrating remarkable robustness against language mismatch and noise. Moreover, given that CASIA provides parallel utterances with identical lexical content across different emotions, the observed improvements can be attributed primarily to paralinguistic cues rather than lexical factors. In particular, the consistent gains of Whisper-based ASR representations over Mel-Spectrogram and SSL models suggest that their advantage stems not only from semantic encoding, but also from a superior ability to capture emotion-relevant prosody, intonation, and spectral dynamics across diverse conditions.
@@ VI Conclusions and Future Work
In this paper, we proposed a novel approach to NSER by leveraging ASR models as robust feature extractors in noisy environments. Our study demonstrated its effectiveness in addressing the limitations of conventional NSER techniques under noise challenges in realistic acoustic environments. By utilizing intermediate-layer representations from ASR models, we captured emotional speech features with greater accuracy, resulting in superior NSER performance compared with conventional noise reduction methods and SSL approaches.
@@
Looking ahead, future work will further build upon the empirical findings of this study by exploring more effective ASR-based NSER frameworks from multiple perspectives. First, given the observed performance variations across different Whisper layers, adaptive or task-aware layer selection strategies could be investigated to more explicitly exploit the complementary roles of encoder and decoder representations. Second, the negative correlation between ASR performance and NSER performance suggests that ASR quality may serve as a useful reference for optimizing downstream emotion recognition under noisy conditions, motivating the design of more tightly coupled ASR–SER co-optimization frameworks. Finally, evaluations will be extended to broader datasets and robustness across different languages and cultural contexts will be explored to further validate the generalization and practical applicability of the proposed framework.